# Approximately-symmetric neural networks for quantum spin liquids

### v1

@@ I Neural quantum states and symmetries of many-body systems
I.1 Overview of different ways of imposing NQS symmetries
Symmetries of many body states ($g|\psi\rangle=|\psi\rangle\ \forall g\in G$) might be imposed on a neural network in multiple ways. If the symmetries of the states are also symmetries of the Hamiltonian, then the approach perhaps most familiar to a physicist is that of mimicking imposing symmetries in the exact diagonalization (ED). As discussed in the main text, in NQS one decomposes an arbitrary state in a certain basis $|\psi\rangle=\sum_{s}\psi_{s}|s\rangle$ where $\{|s\rangle\}$ is typically chosen to be an eigenbasis of the $Z_{j}$ operators. Then a simple way of imposing Hamiltonian symmetries is to turn to the shared eigenbasis of the Hamiltonian and symmetry group $G$. This approach works for an arbitrary group $G$ and has been demonstrated in NQS for e.g., Heisenberg $SU(2)$ symmetric model by [1]. Within such approach one rotates the basis while recalculating matrix elements of the Hamiltonian (in a new basis) and picks a symmetry sector (particular representation of the symmetry group) e.g., $J=0$ for $SU(2)$ by selecting a subset of input bit strings. This approach plays well together with the standard MCMC sampling of bit string configurations in NQS: one initializes the MCMC chain in the symmetry-consistent configuration and then applies a symmetry-preserving update rule. For instance for a $U(1)$ symmetry (e.g., of a quantum $XY$ model) one would only sample configurations with a total spin 00 by initializing the chain accordingly and later using total spin conserving rule for its updates (e.g., one which only exchanges individual spins within the configuration). It should be noted that within this approach the architecture of the neural network does not need to be constrained in any way, yet requires many less parameters to achieve the same relative error in ground state energy (see Fig. 2 in Ref. [1]).
@@ I.1 Overview of different ways of imposing NQS symmetries
I.2 (Approximately) group-equivariant neural networks
@@
Group-equivariant networks [7] are neural network architectures which by construction are equivariant under the action of a particular symmetry group. Group equivariance is one of the essential building blocks for the success of the accurate protein structure prediction with AlphaFold 2 architecture [8]. So far, in the NQS context, group-equivariant networks have mostly been applied within the context of imposing lattice symmetries [9]. Within this approach, one restricts the NQS to (by construction) fulfill a certain constraint: in our case $\psi_{gs}=\psi_{s}$ i.e. that of group-invariance of the output. Here we assume that the group action maps bit strings to bit strings  [^1] Extending the applicability of the method beyond bit string to bit string mapping is perhaps possible by proceeding along lines of Ref. [49] where one imposes $SU(2)$ + lattice symmetries within the constrained neural network architecture utilizing idea of an equivariant Clebsch-Gordan nets [50]..
@@
Critically, generalizing the above approach to approximate symmetries is straightforward: instead of demanding $\psi_{gs}=\psi_{s}\ \ \forall g\in G,\ s\in V$ we demand $\mathbb{E}_{g\in G}$ $\mathbb{E}_{s\sim|\psi(s)|^{2}}||\psi_{gs}-\psi_{s}||<\epsilon$ for some $\epsilon>0$ as discussed in the main text (and where we average over group elements in $G$ and bit strings $s\in V$). We emphasize that $\epsilon$ is learnt by the network itself, perhaps with a help of initialization at the fully symmetric point $\epsilon=0$.
For completeness, we mention that our approach to imposing approximate symmetries within NQS might be perhaps also extended to the framework of Ref. [12]. Therein one evaluates NQS only on “canonical” bit strings which ensures equivariance of the output (for any Abelian group). A “canonical” bit string is a fixed representative of each equivalence class under the group action. Although it is unclear how efficient evaluation of such cannonicalization would be for general Abelian (gauge) groups, we point out an interesting connection to the recent ML literature: [13]. Therein a related cannonicalization function is learnt efficiently by an equivariant neural network. It is thus potentially feasible that the above approximate-symmetries framework might be carried on to this context as well (akin to [14]) - we leave this direction for future work.
I.3 General construction for approximately-symmetric NQS
@@
Let us now discuss how to generalize to the case when the symmetries are only approximate. The basic idea behind any such generalization is to add non-invariant blocks to an otherwise symmetric architecture. We demonstrate this approach with two examples: the “Combo” architecture in the spirit of [21], discussed in the main text, and the residual pathway prior “RPP” architecture in the spirit of [22]. The RPP architecture is constructed to mimic the ResNet [23] skip-connection architecture. Apart from a gauge-invariant pathway, it also contains a non-invariant skip-connection which allows incorporation of non-symmetric features. In practice, for $G=\mathbb{Z}_{2}^{\times N/2}$, we construct the skip connection with a non-$G$-equivariant convolutional layer $\Omega:E\rightarrow P$ which maps set of edges to plaquettes. Finally the two pathways (equivariant and non-equivariant) are summed together and post-processed with another convolutional layer acting on plaquettes (as for the ”Combo” architecture) – see Fig. S1(b) for ”RPP” as compared with ”Combo” in Fig. S1(a).
@@
II Approximately-symmetric NQS performance study
@@ II.1 Simulation parameters
We evaluate the energy $\langle H\rangle$ through Monte Carlo Markov chain (MCMC) sampling. It is performed by noting that $\langle H\rangle=\langle\psi|H|\psi\rangle/\langle\psi|\psi\rangle=\sum_{s}p(s)E_{loc,s}$ where $p(s)=|\psi(s)|^{2}$ and $E_{loc,s}=\sum_{\mathbf{s}^{\prime}}\frac{H_{ss^{\prime}}}{\psi_{s}}\psi_{s}^{\prime}$ where the latter is evaluated exactly for each bit string $s$ (due to only $\mathcal{O}(poly(N))$ non-zero matrix elements of the Hamiltonian). Then for $N_{\textrm{samples}}$ samples one approximates $\langle H\rangle\approx\frac{1}{N_{\textrm{samples}}}\sum_{i=1}^{N_{\textrm{samples}}}E_{loc,s_{i}}$ where samples $\{s_{i}\}$ are found by applying a Metropolis-Hastings algorithm.
We use a custom sampling rule, closely related to the one used in [25], which involves flipping either a single spin per update step, or all the spins surrounding a single vertex. The intution for this update rules stems from the fact that in the ground state of the toric code, a single spin-flip creates (two) excitations, whereas a vertex-flip does not. Thus close to the toric code fixed point, a single-spin flip update would generically be expected to take a high-amplitude state to a low amplitude state, in contrast to a vertex-flip. Nevertheless, single-spin flips are still required for ergodicity. Given $N_{v}$ vertices and $N$ spins, on each update we choose to flip a vertex with probability $p=N_{v}/N$ or else flip a spin. For open boundary conditions, we neglect any verticies with spins on the boundary in this procedure (for periodic boundary conditions, a more complicated procedure is required, see [25]). We find that this vertex-spin update rule empirically yields a significant improvement in MCMC sample acceptance probabilities compared to only using single spin updates.
In order to benefit from GPU parallelization, as well as improving the ergodicity of exploring possibly multi-modal probability landscapes, we draw samples from $\mathcal{O}(10^{3})$ independent MCMC chains processed in parallel. We apply $N_{\textrm{subsample}}$ updates between each collected sample in order to reduce the autocorrelations within each chain. Furthermore, we discard the first $N_{\textrm{burn-in}}$ samples while the chain thermalizes passed its initial transient. We evaluate $\langle H\rangle$ using $C_{L}$ samples per chain, so that the total number of samples is $N_{\textrm{samples}}=C_{L}N_{\textrm{chains}}$. The total number of updates which must be applied is thus $\left(N_{\textrm{burn-in}}+C_{L}\right)N_{\textrm{subsample}}N_{\textrm{chains}}$.
@@
Parameters used for running simulations with the architecture presented in Fig. (1a) of the main text for $16\times 16$ are listed in the Table 1 below. All simulations are run in NetKet [31] that benefits from JAX[32] auto-differentiation and just-in-time compilation. We follow the convention in NetKet 3.10 such that in code neural network outputs represent a logarithm of an amplitude $\psi_{s}$ to facilitate calculations. All neural networks are coded in FLAX [33]. DMRG simulations were performed in iTensor [34, 35] and exact diagonalization in Dynamite [36]. Energy densities as used in the figures are defined as $E_{gs}=\frac{\langle H\rangle}{N+1}$ i.e. energy per stabilizer. Training time in all figures refers to the $\gamma n$ (where $n$ is iteration number, and $\gamma$, as above, is a learning rate).
@@
$N_{\textrm{burn-in}}$ (per chain) $8$
@@
Table 1: Summary of neural network parameters for $16\times 16$ runs. NIB=Non-invariant block, IB=Invariant-block; channels & depth specifies the number of channels in each layer within a block.
II.2 Performance in $h_{x}$ and $h_{z}$ fields for a $4\times 4$ system
@@
II.3 Performance under different perturbations for $4\times 4$
We demonstrate that our ansatz is stable against a variety of different Hamiltonian perturbations: including ones introducing a sign problem. We broadly divide perturbations of the toric code Hamiltonian into different classes (see Table 2). For sign-problem-full perturbations, the Hamiltonian will not be stoquastic, and therefore the Frobenius theorem does not apply, permitting a non-trivial ground-state sign structure. For sign-problem-full perturbations we therefore allow for complex network outputs by using complex parameters.
@@ II.3 Performance under different perturbations for $4\times 4$
Sign-problem-free, invariant XX 1e-9
Sign-problem-free, non-invariant ZZ 6e-8
XX+YY+ZZ 2e-5
XX+YY 1e-5
@@
Table 2: Relative energies obtained within Combo architecture for different perturbations on N=24 toric code. All perturbations are fixed to value $0.2$ in magnitude and in addition we add a small $h_{x}=0.05$ field to all of them.
We note that the performance of our ansatz has accuracy below $10^{-3}$ for all perturbations tested. As generically expected [37], perturbations introducing sign-problem are more difficult for the NQS. This is most likely due to difficulty in propagation of complex phases through the network [38]. We notice that achieving the performance we have reached for the sign-problem-full case hinges upon including correlations between different channels of the non-invariant architecture as depicted in Fig. (1 a) of the main text. For instance, architecture with a ”bottleneck” where one sums over all channels before passing the output to the Wilson loop non-linearity, performs one to two orders of magnitude worse in accuracy as compared with the presented results.
@@
II.4 Performance in $(h_{x},h_{y},h_{z})$ magnetic fields for a $10\times 10$ system
@@
We proceed to investigating what limits the performance of the approximately-symmetric neural network. We first study the performance of the purely gauge-invariant architecture (constructed by omitting non-invariant block in the architecture in Fig. (2a) in the main text). We note that the relative energies (as compared with the state-of-the-art DMRG/QMC results) achieved for $(h_{x},h_{y},h_{z})=(0.2,0.0,0.0)$ fields with a fully gauge-invariant architecture are comparable to these for $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$ with an approximately-symmetric one. We further observe a systematic scaling with the number of parameters and particularly depth of the network. We suggest therefore that the accuracy of the approximately-symmetric network behavior is not limited by the non-invariant block of the network but the invariant one. We investigate this hypothesis by investigating scaling of the relative energy error with the number of channels and depth of the symmetric and non-symmetric block of the network. In agreement with the purely gauge-invariant network results, we observe that by increasing the depth of the invariant part of the network, one can systematically improve the achievable energies (Fig. S3(a)). No similar scaling was observed with the depth of the non-invariant block (Fig. S3(b)) or the number of channels in the non-invariant or invariant part blocks (not shown). This conclusion holds for a range of other values of magnetic fields and system sizes.
Figure S3: Scaling of the relative variational energies with the number of layers (depth) of the neural network architecture on a $10\times 10$ model for (a) the invariant and (b) the non-invariant block under $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$ field. The accuracy of the ansatz can be systematically improved with the depth of the invariant block (but not the non-invariant one). Reference energy corresponds to the depth$=1$ value. (c) Sampling statistics for a $(h_{x},h_{y},h_{z})=(0.2,0.2,0.2)$ on $10\times 10$ system. $\tau$ refers to an energy autocorrelation time. We see that the acceptance rate remains high throughout, while the autocorrelation time does not significantly increase.
@@
We provide more data on MCMC statistics for simulations on a $10\times 10$ system under an $h_{y}$ field to demonstrate that, for the system under consideration and architecture chosen, one does not run into any significant sampling issues. Sampling for the sign-problem full toric code with an $h_{y}\neq 0$ field should be the most difficult: e.g., for stoquastic Hamiltonians it can be proven [42, 43] that MCMC mixing time for sampling their ground states scales polynomially with the number of qubits. To demonstrate that there are no issues with sampling, we show that (Fig. S3(c)): (i) the acceptance probabilities of the MCMC chain remain large ($>25\%$) throughout the optimization; (ii) the energy autocorrelation time $\tau$ is small (where $\tau=0$ is ideal and one would like to avoid $\tau\gg 1$). Furthermore, we verified that the split-$\hat{R}$ characteristics [44] remains $<1.01$ throughout the simulation (where $\hat{R}$ is ideally $1$ and $\hat{R}<1.01$ tolerance is heuristically recommended [44]).
II.5 Extracting phase transitions for a mixed field toric code
@@
For a fixed $h_{z}$ cut through the phase diagram we find no significant shift of the peak in the derivative of both BFFM and second Renyi entropy with increasing $L$ which we associate with the location of the phase transition. For a fixed $h_{x}$ cut, on the other hand, we find stronger finite-size effects  [^2] Lack of symmetry between $X$ and $Z$ perturbations comes from the open boundary conditions: it might be seen in perturbation theory that $X$ and $Z$ magnetic field excitations at the boundary give contributions to energy in different orders of perturbation theory. To obtain a more accurate estimate of the thermodynamic critical points along these cuts, we perform a simple finite-size extrapolation of the position of extracted $\frac{\textrm{d}O_{Z}}{\textrm{d}h_{z}}$ and $\frac{\textrm{d}S_{\textrm{Renyi}}}{\textrm{d}h_{z}}$ peaks (see Fig. S4). In particular, we fit the peak location to a power law in inverse system size, $h_{\textrm{peak}}=h_{\textrm{crit}}+b(\frac{1}{L})^{x}$ and extrapolate $1/L\rightarrow 0$. Uncertainty in the extrapolation determines the uncertainty in critical point extraction (errorbars in Fig. 1(c) in the main text).
@@ III Interpretability
III.1 Approximate to exact symmetries mapping
@@
III.2 Invariance error
In the definition of the approximate group invariance $\mathbb{E}_{s\sim p(s)}\mathbb{E}_{g\in G}|\psi_{gs}-\psi_{s}|<\epsilon$, we defined a measure of the magnitude of the symmetry breaking, $\epsilon$. Here we discuss the dynamics and asymptotic values of the closely related quantity $\tilde{\epsilon}=\mathbb{E}_{s\sim p(s)}\mathbb{E}_{g\in G}\|\log\psi_{gs}-\log{\psi_{s}}\|$ ) ∥ as learnt by the neural network. We note that: (i) value of $\tilde{\epsilon}$ approaches the same value irrespective of the network hyperparameters over the course of optimization, (ii) $\tilde{\epsilon}$ appears to have a divergent derivative as one tunes the symmetry-violating field in a point closely matching the more conventional observables of the toric code phase transitions such as BFFM order parameters (see main text).

### v2

@@ Abstract
S1.1 Overview of different ways of imposing NQS symmetries
Symmetries of many body states ($g|\psi\rangle=|\psi\rangle\ \forall g\in G$) might be imposed on a neural network in multiple ways. If the symmetries of the states are also symmetries of the Hamiltonian, then the approach perhaps most familiar to a physicist is that of mimicking imposing symmetries in the exact diagonalization (ED). As discussed in the main text, in NQS one decomposes an arbitrary state in a certain basis $|\psi\rangle=\sum_{s}\psi_{s}|s\rangle$ where $\{|s\rangle\}$ is typically chosen to be an eigenbasis of the $Z_{j}$ operators. Then a simple way of imposing Hamiltonian symmetries is to turn to the shared eigenbasis of the Hamiltonian and symmetry group $G$. This approach works for an arbitrary group $G$ and has been demonstrated in NQS for e.g., Heisenberg $SU(2)$ symmetric model by [71]. Within such approach one rotates the basis while recalculating matrix elements of the Hamiltonian (in a new basis) and picks a symmetry sector (particular representation of the symmetry group) e.g., $J=0$ for $SU(2)$ by selecting a subset of input bit strings. This approach plays well together with the standard Monte Carlo Markov Chain (MCMC) sampling of bit string configurations in NQS: one initializes the MCMC chain in the symmetry-consistent configuration and then applies a symmetry-preserving update rule. For instance for a $U(1)$ symmetry (e.g., of a quantum $XY$ model) one would only sample configurations with a total spin 0 by initializing the chain accordingly and later using total spin conserving rule for its updates (e.g., one which only exchanges individual spins within the configuration). It should be noted that within this approach the architecture of the neural network does not need to be constrained in any way, yet requires many less parameters to achieve the same relative error in ground state energy (see Fig. 2 in Ref. [71]).
@@
S1.2 (Approximately) group-equivariant neural networks
@@
Group-equivariant networks [36] are neural network architectures which by construction are equivariant under the action of a particular symmetry group. Group equivariance is one of the essential building blocks for the success of the accurate protein structure prediction with AlphaFold 2 architecture [74]. So far, in the NQS context, group-equivariant networks have mostly been applied within the context of imposing lattice symmetries [30]. Within this approach, one restricts the NQS to (by construction) fulfill a certain constraint: in our case $\psi_{gs}=\psi_{s}$ i.e. that of group-invariance of the output. Here we assume that the group action maps bit strings to bit strings. (Extending the applicability of the method beyond bit string to bit string mapping is perhaps possible by proceeding along lines of Ref. [70] where one imposes $SU(2)$ + lattice symmetries within the constrained neural network architecture utilizing idea of an equivariant Clebsch-Gordan nets [75]).
@@
[…]
For completeness, we mention that our approach to imposing approximate symmetries within NQS might be perhaps also extended to the framework of Ref. [77]. Therein one evaluates NQS only on “canonical” bit strings which ensures equivariance of the output (for any Abelian group). A “canonical” bit string is a fixed representative of each equivalence class under the group action. Although it is unclear how efficient evaluation of such canonicalization would be for general Abelian (gauge) groups, we point out an interesting connection to the recent ML literature: [78]. Therein a related canonicalization function is learnt efficiently by an equivariant neural network. It is thus potentially feasible that the above approximate-symmetries framework might be carried on to this context as well (akin to [79]) - we leave this direction for future work.
S1.3 General construction for approximately-symmetric NQS
@@
Let us now discuss how to generalize to the case when the symmetries are only approximate. The basic idea behind any such generalization is to add non-invariant blocks to an otherwise symmetric architecture. We demonstrate this approach with two examples: the “Combo” architecture in the spirit of [43], discussed in the main text, and the residual pathway prior “RPP” architecture in the spirit of [42]. The RPP architecture is constructed to mimic the ResNet [85] skip-connection architecture. Apart from a gauge-invariant pathway, it also contains a non-invariant skip-connection which allows incorporation of non-symmetric features. In practice, for $G=\mathbb{Z}_{2}^{\times N/2}$, we construct the skip connection with a non-$G$-equivariant convolutional layer $\Omega:E\rightarrow P$ which maps set of edges to plaquettes. Finally the two pathways (equivariant and non-equivariant) are summed together and post-processed with another convolutional layer acting on plaquettes (as for the “Combo” architecture) – see Fig. S1(b) for “RPP” as compared with “Combo” in Fig. S1(a).
@@
S1.4 Potential use cases for approximately-symmetric models
We emphasize that there exists many target systems with multiple open physics questions where our approximately-symmetric neural network may be applied. These can be addressed using our approximately symmetric architecture, with minor extensions, for both ground state and real-time evolution studies. For clarity we summarize them in the Table S1.
Extension Motivation Framework modification
$\mathds{Z}_{2}$ Spin Liquids Realizable in quantum simulators via engineering [5, 4]; resulting ground states (and their dynamical preparation) away from exactly symmetric fixed point. No modification
Chiral $\mathds{Z}_{d}$ toric code $\mathds{Z}_{d}$ has an exact symmetry limit and for $d>2$ permits chiral (exact local symmetry-breaking) perturbations; allows study of fractional quantum Hall physics [86]. Increase onsite dimension $d$, modify group nonlinearity [38]: $x_{1}x_{2}x_{3}x_{4}\rightarrow x_{1}x_{2}x^{*}_{3}x^{*}_{4}$ (around each plaquette)
$U(1)$ lattice gauge theory Approximately realizable in quantum simulators via engineering [87]; study fractionalization and confinement dynamics in $U(1)$ gauge theories Similar to $\mathds{Z}_{d}$; promote input variables from binary to continuous [39].
Kitaev double Study phases and dynamics of models with discrete non-abelian symmetries and non-abelian anyons [7]. Similar to $\mathds{Z}_{d}$ with special care taken for multiplication ordering [38].
Double semion Study of sign-problem-full [88] twisted gauge theories featuring semionic excitations [89]. Modify product non-linearity from group invariance to group equivariance [77, 78].
Non-abelian local continuous symmetries Study of realizations of non-Abelian lattice gauge theories such as QCD in quantum simulators [90]. Generalizing framework to non-Abelian continuous symmetries [69].
Approximate global symmetries Study of weakly positionally disordered systems e.g., with approximate translation symmetry. Using group equivariant neural networks with relaxed group convolutions [43].
Table S1: Partial list of models where approximately symmetric neural networks may be helpful and the extensions that they would require.
Appendix S2 Approximately-symmetric NQS toric code performance study
@@
We evaluate the energy $\langle H\rangle$ through MCMC sampling. It is performed by noting that $\langle H\rangle=\langle\psi|H|\psi\rangle/\langle\psi|\psi\rangle=\sum_{s}p(s)E_{loc,s}$ where $p(s)=|\psi(s)|^{2}$ and $E_{loc,s}=\sum_{\mathbf{s}^{\prime}}\frac{H_{ss^{\prime}}}{\psi_{s}}\psi_{s}^{\prime}$ where the latter is evaluated exactly for each bit string $s$ (due to only $\mathcal{O}(poly(N))$ non-zero matrix elements of the Hamiltonian). Then for $N_{\textrm{samples}}$ samples one approximates $\langle H\rangle\approx\frac{1}{N_{\textrm{samples}}}\sum_{i=1}^{N_{\textrm{samples}}}E_{loc,s_{i}}$ where samples $\{s_{i}\}$ are found by applying a Metropolis-Hastings algorithm.
We use a custom sampling rule, closely related to the one used in [92], which involves flipping either a single spin per update step, or all the spins surrounding a single vertex. The intuition for this update rule stems from the fact that in the ground state of the toric code, a single spin-flip creates (two) excitations, whereas a vertex-flip does not. Thus close to the toric code fixed point, a single-spin flip update would generically be expected to take a high-amplitude state to a low amplitude state, in contrast to a vertex-flip. Nevertheless, single-spin flips are still required for ergodicity. Given $N_{v}$ vertices and $N$ spins, on each update we choose to flip a vertex with probability $p=N_{v}/N$ or else flip a spin. For open boundary conditions, we neglect any vertex with spins on the boundary in this procedure (for periodic boundary conditions, a more complicated procedure is required, see [92]). We find that this vertex-spin update rule empirically yields a significant improvement in MCMC sample acceptance probabilities compared to only using single spin updates.
In order to benefit from GPU parallelization, as well as improving the ergodicity of exploring possibly multi-modal probability landscapes, we draw samples from $\mathcal{O}(10^{3})$ independent MCMC chains processed in parallel. We apply $N_{\textrm{subsample}}$ updates between each collected sample in order to reduce the autocorrelations within each chain. Furthermore, we discard the first $N_{\textrm{burn-in}}$ samples while the chain thermalizes past its initial transient. We evaluate $\langle H\rangle$ using $C_{L}$ samples per chain, so that the total number of samples is $N_{\textrm{samples}}=C_{L}N_{\textrm{chains}}$. The total number of updates which must be applied is thus $\left(N_{\textrm{burn-in}}+C_{L}\right)N_{\textrm{subsample}}N_{\textrm{chains}}$.
@@
Parameters used for running simulations with the architecture presented in Fig. (1)a of the main text for $16\times 16$ are listed in the Table S2 below. All simulations are run in NetKet [52] that benefits from JAX[95] auto-differentiation and just-in-time compilation. We follow the convention in NetKet 3.16 such that in code neural network outputs represent a logarithm of an amplitude $\psi_{s}$ to facilitate calculations. All neural networks are coded in FLAX [96]. Density Matrix Renormalization Group (DMRG) simulations were performed in iTensor [97, 98] and exact diagonalization in Dynamite [99]. Energy densities as used in the figures are defined as $E_{gs}=\frac{\langle H\rangle}{N+1}$ i.e. energy per stabilizer. Training time in all figures refers to the $\gamma n$ (where $n$ is iteration number, and $\gamma$, as above, is a learning rate).
@@
$N_{\textrm{burn-in}}$ (per chain) 88
@@
Table S2: Summary of neural network parameters for $16\times 16$ runs. NIB=Non-invariant block, IB=Invariant-block; channels & depth specifies the number of channels in each layer within a block.
S2.2 Comparisons with baseline neural networks for a $4\times 4$ system
We provide more details on comparing our approach with baseline neural network architectures which incorporate lattice symmetries but no local gauge symmetries. In the main text this was performed for CNNs, which by construction incorporate translation symmetries and contain 4104-10 layers with multiple filters per layer. We note that the main difference between CNNs and our architecture is the presence of the group-invariant block and both networks have a comparable runtime. As mentioned in the main text (cf. Fig. 2(a) therein), we found that unlike our approximately-symmetric network approach, CNNs are trapped in local minima, thus conclusively showing the importance of imposing approximate gauge symmetries. Additionally, in Fig. S2, we show that symmetric restricted Boltzmann machines (RBMs), which incorporate all lattice symmetries rather than only translations [21], perform better than CNNs. However, they still get trapped in local minima with errors several orders of magnitude larger than those achieved by our architecture. Furthermore, we expect the performance gap between symmetric RBMs (or their deeper variants, such as GCNNs) and approximate symmetry-based approaches to grow with system size, as the local symmetry space scales exponentially ($\mathcal{O}(2^{N/2})$) compared to the linear scaling of lattice symmetries ($\mathcal{O}(N)$).
Figure S2: Energy convergence curve for approximately-symmetric networks and RBMs as compared with exact diagonalization on $N=24$ system in $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$ fields.
S2.3 Performance in $h_{x}$ and $h_{z}$ fields for a $4\times 4$ system
@@
S2.4 Performance under different perturbations for $4\times 4$
We demonstrate that our ansatz is stable against a variety of different Hamiltonian perturbations: including ones introducing a sign problem. We broadly divide perturbations of the toric code Hamiltonian into different classes (see Table S3). For sign-problem-full perturbations, the Hamiltonian will not be stoquastic, and therefore the Frobenius theorem does not apply, permitting a non-trivial ground-state sign structure. For sign-problem-full perturbations we therefore allow for complex network outputs by using complex parameters.
@@
Sign-problem-free, invariant $X$ 1e-9
Sign-problem-free, non-invariant $Z$ 6e-8
XX+YY+$Z$ 2e-5
XX+$Y$ 1e-5
@@
Table S3: Relative energies obtained within Combo architecture for different perturbations on N=24 toric code. All perturbations are fixed to a value of $0.2$ in magnitude and in addition we add a small $h_{x}=0.05$ field to all of them.
We note that the performance of our ansatz has accuracy below $10^{-3}$ for all perturbations tested. As generically expected [100], perturbations introducing sign-problem are more difficult for the NQS. This is most likely due to difficulty in propagation of complex phases through the network [101]. We notice that achieving the performance we have reached for the sign-problem-full case hinges upon including correlations between different channels of the non-invariant architecture as depicted in Fig. 1(a) of the main text. For instance, architecture with a ”bottleneck” where one sums over all channels before passing the output to the Wilson loop non-linearity, performs one to two orders of magnitude worse in accuracy as compared with the presented results.
@@
S2.5 Performance in $(h_{x},h_{y},h_{z})$ magnetic fields for a $10\times 10$ system
@@
We proceed to investigating what limits the performance of the approximately-symmetric neural network. We first study the performance of the purely gauge-invariant architecture (constructed by omitting non-invariant block in the architecture in Fig. (2a) in the main text). We performed additional validations to characterize the converged ground states—such as evaluating order parameters for the topological phase—and found that the V-score [103] was of the same order of magnitude as the relative error compared to state-of-the-art DMRG/QMC results, consistent with expectations. Additionally, we note that the relative energies achieved for $(h_{x},h_{y},h_{z})=(0.2,0.0,0.0)$ fields with a fully gauge-invariant architecture are comparable to these for $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$ with an approximately-symmetric one. We further observe a systematic scaling with the number of parameters and particularly depth of the network. We suggest therefore that the accuracy of the approximately-symmetric network behavior is not limited by the non-invariant block of the network but the invariant one. We investigate this hypothesis by investigating scaling of the relative energy error with the number of channels and depth of the symmetric and non-symmetric block of the network. In agreement with the purely gauge-invariant network results, we observe that by increasing the depth of the invariant part of the network, one can systematically improve the achievable energies (Fig. S4(a)). No similar scaling was observed with the depth of the non-invariant block (Fig. S4(b)) or the number of channels in the non-invariant or invariant part blocks (not shown). This conclusion holds for a range of other values of magnetic fields and system sizes.
Figure S4: Scaling of the relative variational energies with the number of layers (depth) of the neural network architecture on a $10\times 10$ model for (a) the invariant and (b) the non-invariant block under $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$ field. The accuracy of the ansatz can be systematically improved with the depth of the invariant block (but not the non-invariant one). Reference energy corresponds to the depth=1=1 value. (c) Sampling statistics for a $(h_{x},h_{y},h_{z})=(0.2,0.2,0.2)$ on $10\times 10$ system. $\tau$ refers to an energy autocorrelation time. We see that the acceptance rate remains high throughout, while the autocorrelation time does not significantly increase.
@@
We provide more data on MCMC statistics for simulations on a $10\times 10$ system under an $h_{y}$ field to demonstrate that, for the system under consideration and architecture chosen, one does not run into any significant sampling issues. Sampling for the sign-problem full toric code with an $h_{y}\neq 0$ field should be the most difficult: e.g., for stoquastic Hamiltonians it can be proven [106, 107] that MCMC mixing time for sampling their ground states scales polynomially with the number of qubits. To demonstrate that there are no issues with sampling, we show that (Fig. S4(c)): (i) the acceptance probabilities of the MCMC chain remain large ($25\%$) throughout the optimization; (ii) the energy autocorrelation time $\tau$ is small (where $\tau=0$ is ideal and one would like to avoid $\tau\gg 1$). Furthermore, we verified that the split-$\hat{R}$ characteristics [108] remains <1.01<1.01 throughout the simulation (where $\hat{R}$ is ideally $1$ and $\hat{R}<1.01$ tolerance is heuristically recommended [108]).
S2.6 Extracting phase transitions for a mixed field toric code
@@
For a fixed $h_{z}$ cut through the phase diagram we find no significant shift of the peak in the derivative of both BFFM and second Renyi entropy with increasing $L$ which we associate with the location of the phase transition. For a fixed $h_{x}$ cut, on the other hand, we find stronger finite-size effects. (Lack of symmetry between $X$ and $Z$ perturbations comes from the open boundary conditions: it might be seen in perturbation theory that $X$ and $Z$ magnetic field excitations at the boundary give contributions to energy in different orders of perturbation theory). To obtain a more accurate estimate of the thermodynamic critical points along these cuts, we perform a simple finite-size extrapolation of the position of extracted $\frac{\textrm{d}O_{Z}}{\textrm{d}h_{z}}$ and $\frac{\textrm{d}S_{\textrm{Renyi}}}{\textrm{d}h_{z}}$ peaks (see Fig. S5). In particular, we fit the peak location to a power law in inverse system size, $h_{\textrm{peak}}=h_{\textrm{crit}}+b(\frac{1}{L})^{x}$ and extrapolate $1/L\rightarrow 0$. Uncertainty in the extrapolation determines the uncertainty in critical point extraction (errorbars in Fig. 1(c) in the main text).
@@
S3.1 Approximate to exact symmetries mapping
@@
S3.2 Invariance error
[…]